# The adjoint state method for parametric definable optimization without smoothness or uniqueness

### 27 Mar 2026 v1

@@ 2.3 Comparison to alternative methods
In most nonlinear programming solvers, such as SNOPT [25] and IPOPT [40], dual variables are already computed internally as part of satisfying the KKT optimality conditions, or because of the primal-dual nature of the underlying algorithms. Consequently, in such a situation, getting the dual solution in addition to the primal solution, requires little to no additional computational effort, $\text{Cost}\big(\mathrm{solver_{pd}}\big)$ is the same as $\text{Cost}\big(\mathrm{solver_{p}}\big)$ and the adjoint state formula ASM is readily implementable with minimal code extension.
@@
$\text{Cost}\big(\text{AD}\big)$ $=\text{Cost}\big(\mathrm{solver_{p}}\big)+\text{Cost}\big(\texttt{autodiff}F(\mathrm{solver_{p}}(\bar{\theta}),\bar{\theta})\big)+\text{Cost}\big(\texttt{autodiff}\,\mathrm{solver_{p}}(\bar{\theta})\big)$
@@ 3.2 Definition of o-minimality
$\bigcap_{k=1}^{K}\{x\in\mathbb{R}^{\ell}:\phi_{k}(x)<0,\;\varphi_{k}(x)=0\},$
where $K\geq 1$ and $\phi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and $\varphi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ are polynomial functions for all $k\in\{1,\ldots,K\}$. Therefore, the semialgebraic functions and semialgebraic set-valued maps are functions and set-valued maps, respectively, whose graphs are semialgebraic sets.
@@ 4 The adjoint state method: general conservative case
4.1 Preliminaries
@@
$R(\theta)\coloneq\left\{x\in\mathbb{R}^{n}:G(x,\theta)\leq 0,H(x,\theta)=0\right\},$
@@
Remark 4.1.
By continuity of $G$ and $H$, the set-valued map $R$ is graph closed.
Our main construction is based on the following.
@@
    The objective function $F$ is definable with definable conservative field $D_{F}$. The constraint functions $G$ and $H$ are also definable. In that case, we call Problem (2.1) a parametric definable program.
@@
    The MFCQ is satisfied at any $(\theta,x)\in\operatorname*{gph}R$, i.e., the MFCQ holds throughout the feasible set $R(\theta)$ for any $\theta\in\mathbb{R}^{d}$.
@@
[…]
@@
[…]
We set $D_{F(\cdot,\theta)}\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$, the projection on the $x$ component of $D_{F}$, which provides a conservative for the partial function $F(\cdot,\theta)$, for fixed $\theta\in\mathbb{R}^{q}$.
Let $\operatorname*{Mult}\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{m}\times\mathbb{R}^{p}$ be the set-valued map providing Lagrange multipliers satisfying stationarity and complementarity from the KKT system
$(x,\theta)$ $\mapsto\Big\{(\lambda,\mu)\in\mathbb{R}^{m}\times\mathbb{R}^{p}:0\in\operatorname*{conv\!}D_{F(\cdot,\theta)}(x)+\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,$
$\qquad\quad\lambda\geq 0,\lambda^{\!\mathsf{T}}G(x,\theta)=0\Big\},$
We then define the set-valued map of primal-dual solutions $\mathrm{Sol_{pd}}\colon\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ as
$\mathrm{Sol_{pd}}(\theta)\coloneq\{(x,\lambda,\mu)\in\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}:x\in\mathrm{Sol_{p}}(\theta),(\lambda,\mu)\in\operatorname*{Mult}(x,\theta)\}.$
@@
For all $\theta\in\mathbb{R}^{q}$ and for every $x\in\mathrm{Sol_{p}}(\theta)$, we have the conservative KKT conditions given by [29, Theorem 3.3] and [8, Corollary 1]. Then $\operatorname*{Mult}(\theta,x)\neq\varnothing$ and $\mathrm{Sol_{pd}}(\theta)$ is nonempty.
Let $\theta\in\mathbb{R}^{q}$. For the sake of contradiction, let’s assume that $\mathrm{Sol_{pd}}$ is not locally bounded around $\theta$. It follows that there exist $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ and $(x_{k},\lambda_{k},\mu_{k})\in\mathrm{Sol_{pd}}(\theta_{k})$, for all $k\in\mathbb{N}$, such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta$ and $\lVert{(x_{k},\lambda_{k},\mu_{k})}\rVert\xrightarrow[k\to\infty$$\infty$. Since $\mathrm{Sol_{p}}$ is locally bounded, we have that $(x_{k})_{k\in\mathbb{N}}$ is bounded and $\lVert{(\lambda_{k},\mu_{k})}\rVert\xrightarrow[k\to\infty$$\infty$. Passing to subsequences if necessary, we can assume that $\lVert{(\lambda_{k},\mu_{k})}\rVert\neq 0$ for all $k\in\mathbb{N}$, and $x_{k}\xrightarrow[k\to\infty]{}x$. By Lemma A.2(), we get $x\in\mathrm{Sol_{p}}(\theta)$. Let, for all $k\in\mathbb{N}$, $\alpha_{k}=\frac{\lambda_{k}}{\lVert{(\lambda_{k},\mu_{k})}\rVert}$ and $\beta_{k}=\frac{\mu_{k}}{\lVert{(\lambda_{k},\mu_{k})}\rVert}$. So for any $k\in\mathbb{N}$, $\lVert{\alpha_{k}}\rVert\leq 1$ and $\lVert{\beta_{k}}\rVert\leq 1$, and, without loss of generality, we can also assume $\alpha_{k}\xrightarrow[k\to\infty]{}\alpha$ and $\beta_{k}\xrightarrow[k\to\infty]{}\beta$. For all $k\in\mathbb{N}$, we have that $\lVert{(\alpha_{k},\beta_{k})}\rVert=1$. It is clear that $\lVert{(\alpha,\beta)}\rVert=1$. By definition of the set-valued map $\operatorname*{Mult}$, we know that, for every $k\in\mathbb{N}$,
$\exists\,v_{k}\in\operatorname*{conv\!}D_{F(\cdot,\theta_{k})}(x_{k}),\operatorname{Jac}_{x}\!G(x_{k},\theta_{k})^{\!\mathsf{T}}\alpha_{k}+\operatorname{Jac}_{x}\!H(x_{k},\theta_{k})^{\!\mathsf{T}}\beta_{k}=-\frac{v_{k}}{\lVert{(\alpha_{k},\beta_{k})}\rVert}.$ (4.1)
Using the fact that $(x,\theta)\mapsto D_{F(\cdot,\theta)}(x)$ is locally bounded (because $D_{F}$ is locally bounded), we have that $(v_{k})_{k\in\mathbb{N}}$ is bounded. Again, passing to a subsequence if needed, we have that $v_{k}\xrightarrow[k\to\infty]{}v$. Taking the limit in (4.1), we can write
$\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\alpha+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\beta=0.$ (4.2)
Furthermore, for all $k\in\mathbb{N}$ and $i\in\{1,\ldots,m\}$,
$\alpha_{k,i}\geq 0\;\text{ and }\;\alpha_{k,i}g_{i}(x_{k},\theta_{k})=0.$
Passing to the limit, we get, for all $i\in\{1,\ldots,m\}$,
$\alpha_{i}\geq 0\;\text{ and }\;\alpha_{i}g_{i}(x,\theta)=0.$ (4.3)
Given Equations (4.2) and (4.3), the dual form of the MFCQ in (2.1) imposes $\alpha=\beta=0$. This contradicts the fact that $\lVert{(\alpha,\beta)}\rVert=1$. Thus $\mathrm{Sol_{pd}}$ is locally bounded around $\theta$. ∎
@@
The last equality in (4.4) is due to the fact that, for all $(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta)$, one has by complementarity
$x\in\mathrm{Sol_{p}}(\theta),\;\lambda^{\!\mathsf{T}}G(x,\theta)=0,\text{ and }H(x,\theta)=0.$
@@ 4.2 Parametric subdifferentiation of the value function
As discussed earlier, [36, Theorem 10.13] provides a formula for the subdifferential of $f$, which has a strong connection with the adjoint formula in (1.2). Let us make this explicit in the context of Problem (1.1). We will consider, only in this section, the extended formulation as follows
@@
$(x,\theta)$ $\quad\mapsto\quad F(x,\theta)+\delta_{\mathbb{R}^{m}_{-}}(G(x,\theta))+\delta_{\{0^{p}\}}(H(x,\theta)),$
where, for $C\in\mathbb{R}^{\ell}$ with $\ell\in\mathbb{N}$, $\delta$ is the indicator function of $C$, i.e., $\delta_{C}(y)=0$ if $y\in C$ and $\delta_{C}(y)=\infty$ otherwise, and $\{0^{p}\}$ denotes the origin in $\mathbb{R}^{p}$. Problem (1.1) is equivalent to unconstrained partial minimization of $\bar{F}$ so that this fits the context of [36, Theorem 10.13].
Under Assumption 4.2 the constraint qualification holds jointly in $x,\theta$, see Remark 2.2 and the chain rule [36, Theorem 10.6] applies to the indicators in the definition of $\bar{F}$. Combining this with the sum rule [36, Exercise 10.10], we have, for all $(x,\theta)\in\operatorname*{gph}R$,
$\partial^{c}\bar{F}(x,\theta)$ $\subset\partial^{c}F(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta),\operatorname{Jac}_{\theta}\!G(x,\theta)\right)^{\!\mathsf{T}}N_{\mathbb{R}^{m}_{-}}(G(x,\theta))$
$\qquad+\left(\operatorname{Jac}_{x}\!H(x,\theta),\operatorname{Jac}_{\theta}\!H(x,\theta)\right)^{\!\mathsf{T}}N_{\{0^{p}\}}(H(x,\theta)),$
where, for $C\subset\mathbb{R}^{\ell}$ convex with $\ell\in\mathbb{N}$ and $y\in\mathbb{R}^{\ell}$, $N_{C}(y)$ is the normal cone of $C$ at $y$. In our case the elements of the normal cones are given by the Lagrange multipliers $\lambda,\mu$ in $\mathrm{Sol_{pd}}(\theta)$. Putting everything together, the parametric optimality condition, [36, Theorem 10.13], ensures that, under Assumptions 4.2(), for all $\theta\in\mathbb{R}^{q}$,
$\partial^{c}f(\theta)\subset D_{f}(\theta)\coloneq\operatorname*{conv\!}\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),$ (4.5)
$(0,u)\in\partial^{c}F(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}.$
Importantly, for all $\theta\in\mathbb{R}^{q}$ and for all $x\in\mathrm{Sol_{p}}(\theta)$, thanks to the parametric Fermat’s rule [36, Example 10.12], there exists $u\in\mathbb{R}^{q}$ such that $(0,u)\in\partial^{c}\bar{F}(x,\theta)$; hence $D_{f}(\theta)$ is nonempty.
However, as already stated earlier, the inclusion in (4.5) is not sharp in general and the generalized adjoint formula produces artifacts, see Example 2.7. We show in the coming section that, under Assumptions 4.2(), the set-valued map $D_{f}$ is a conservative field for $f$, providing a conservative adjoint formula.

### 22 Apr 2026 v2

@@ 2.3 Algorithmic differentiation and alternative methods
In most nonlinear programming solvers, such as SNOPT [29] and IPOPT [47], dual variables are already computed internally as part of satisfying the KKT optimality conditions, or because of the primal-dual nature of the underlying algorithms. This also applies to conic programming solvers such as SDPA [49], SDPT3 [45] or commercial solvers such as MOSEK or KNITRO. Consequently, in such a situation, getting the dual solution in addition to the primal solution, requires little to no additional computational effort, $\text{Cost}\big(\mathrm{solver_{pd}}\big)$ is the same as $\text{Cost}\big(\mathrm{solver_{p}}\big)$ and the adjoint state formula ASM() is readily implementable with minimal code extension.
@@
$\text{Cost}\big(\text{AD}\big)$ $=\text{Cost}\big(\mathrm{solver_{p}}\big)+\text{Cost}\big(\texttt{autodiff}F(\mathrm{solver_{p}}(\bar{\theta}),\bar{\theta})\big)$
$\qquad+\text{Cost}\big(\texttt{autodiff}\,\mathrm{solver_{p}}(\bar{\theta})\big)$
@@ 3.2 Definition of o-minimality
$\bigcap_{k=1}^{\bar{k}}\{x\in\mathbb{R}^{\ell}:\phi_{k}(x)<0,\;\varphi_{k}(x)=0\},$
where $\bar{k}\geq 1$ and $\phi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ and $\varphi_{k}\colon\mathbb{R}^{\ell}\to\mathbb{R}$ are polynomial functions for all $k\in\{1,\ldots,\bar{k}\}$. Therefore, the semialgebraic functions and semialgebraic set-valued maps are functions and set-valued maps, respectively, whose graphs are semialgebraic sets.
@@ 4 The adjoint state method: general conservative case
In this section, we revisit the general problem (1.1), which has the following form: for a closed convex cone $K\subset\mathbb{R}^{m}$, a locally Lipschitz function $F\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}$, and a continuously differentiable mapping $C\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}$,
$f(\theta)\coloneq\inf_{x\in\mathbb{R}^{n}}\left\{F(x,\theta)\;:\;C(x,\theta)\in K\right\}.$ (4.1)
Throughout the remainder of this section, all notation is adapted to Problem (4.1). In this setting, the MFCQ condition given in Equation (2.1) of Definition 2.1() extends as follows. We denote by $K^{\circ}\coloneq\left\{x\in\mathbb{R}^{m}\;:\;\langle x,y\rangle\leq 0,\ \forall y\in K\right\}$ the polar cone of $K$.
Definition 4.1 (RCQ).
The Robinson’s Constraint Qualification is satisfied at $(x,\theta)\in\mathbb{R}^{n}\times\mathbb{R}^{q}$ whenever
$\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\lambda=0,\,\lambda\in K^{\circ},\,\left\langle C(x,\theta),\lambda\right\rangle=0\implies\lambda=0.$
The RCQ in Definition 4.1() is one of the equivalent forms of Robinson’s constraint qualification in finite dimension [11, Corollary 2.98]. Note that the condition on $\lambda$ expresses the fact that $\lambda\in N_{K}(C(x,\theta))$, where $N_{K}$ denotes the normal cone to $K$ as follows.
Lemma 4.2.
([43, Corollary 23.5.4]) Let $K\subset\mathbb{R}^{m}$ be a closed convex cone, then for any $z\in K$, we have $N_{K}(z)=\{\lambda\in K^{\circ},\,\left\langle z,\lambda\right\rangle=0\}$.
The condition $\left\langle z,\lambda\right\rangle=0$ in Lemma 4.2() is called complementarity. It is a specificity of convex cones and constitutes a crucial proof element. The Lagrangian of Problem (4.1), $\mathcal{L}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{q}\to\mathbb{R}$, is the same as for (2.1) and writes
$\mathcal{L}(x,\lambda,\theta)\coloneq F(x,\theta)+\lambda^{\!\mathsf{T}}C(x,\theta).$
We recover (2.1) using the following lemma.
Lemma 4.3.
Set $K=\mathbb{R}^{m}_{-}\times\{0\}\subset\mathbb{R}^{m}\times\mathbb{R}^{p}$ and let $C\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\to\mathbb{R}^{m}\times\mathbb{R}^{p}$ be continuously differentiable and defined blockwise with first component $G$ and second component $H$, then the MFCQ condition from Definition 2.1() holds if and only if the RCQ condition in Definition 4.1() holds.
It is routine to check that (2.1) is the same as RCQ using the fact that $K^{\circ}=\mathbb{R}^{m}_{+}\times\mathbb{R}^{p}$. ∎
4.1 Preliminary results
@@
$R(\theta)\coloneq\left\{x\in\mathbb{R}^{n}:C(x,\theta)\in K\right\},$
@@
Note that, by continuity of $C$ and closedness of $K$, the set-valued map $R$ is graph closed. Our main construction is based on the following.
@@
    The objective function $F$ is locally Lipschitz and definable with definable conservative field $D_{F}$. The constraint function $C$ is continuously differentiable and definable, and $K$ is a closed convex definable cone.
@@
    The RCQ is satisfied at any $(\theta,x)\in\operatorname*{gph}R$, i.e., the RCQ from Definition 4.1() holds throughout the feasible set $R(\theta)$ for any $\theta\in\mathbb{R}^{q}$.
@@
    When $K$ is the negative orthant, the authors in [6] proved that, for diagonally perturbed definable programming, the Assumption $4\tagform@ii$() is generic, i.e., the MFCQ holds throughout the feasible set for all but finitely many (hence for Lebesgue-almost all) diagonal perturbations; Problem (2.1) with no equality constraints corresponds to the zero diagonal perturbation. This result hints at the prevalence of the MFCQ in the definable context.
@@
    Assumption $4\tagform@iii$() ensures that, for any $\theta\in\mathbb{R}^{q}$, the infimum in the definition of the value function $f$, in (4.1), is achieved. Then $f$ can be equivalently defined as $\theta\mapsto F(\mathrm{Sol_{p}}(\theta),\theta)$.
We set $D_{F(\cdot,\theta)}\colon\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$, the projection on the $x$ component of $D_{F}$, which provides a conservative fields for the partial function $F(\cdot,\theta)$, for fixed $\theta\in\mathbb{R}^{q}$. The KKT system for Problem (4.1) is given by
$0\in\operatorname*{conv\!}D_{F(\cdot,\theta)}(x)+\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\lambda,\quad C(x,\theta)\in K,\,\lambda\in K^{\circ},\,\left\langle C(x,\theta),\lambda\right\rangle=0,$ (4.2)
where the conditions on $\lambda$ express the fact that $\lambda\in N_{K}(C(x,\theta))$ as in Lemma 4.2(). Let $\operatorname*{Mult}\colon\mathbb{R}^{n}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{m}$ be the set-valued map providing Lagrange multipliers satisfying stationarity and complementarity from the KKT system (4.2):
$\operatorname*{Mult}(x,\theta)\coloneq\big\{\lambda\in\mathbb{R}^{m}:0\in\operatorname*{conv\!}D_{F(\cdot,\theta)}(x)+\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\lambda,\,\,\lambda\in K^{\circ},\lambda^{\!\mathsf{T}}C(x,\theta)=0\big\}.$
We then define the set-valued map of primal-dual solutions $\mathrm{Sol_{pd}}\colon\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{m}$ as
$\mathrm{Sol_{pd}}(\theta)\coloneq\{(x,\lambda)\in\mathbb{R}^{n}\times\mathbb{R}^{m}:x\in\mathrm{Sol_{p}}(\theta),\lambda\in\operatorname*{Mult}(x,\theta)\}.$
@@
For all $\theta\in\mathbb{R}^{q}$ and for every $x\in\mathrm{Sol_{p}}(\theta)$, the conservative KKT conditions (4.2) hold true by combining [34, Theorem 3.3] and [9, Corollary 1], so that $\operatorname*{Mult}(x,\theta)\neq\varnothing$ and $\mathrm{Sol_{pd}}(\theta)$ is nonempty.
Let $\theta\in\mathbb{R}^{q}$. For the sake of contradiction, we assume that $\mathrm{Sol_{pd}}$ is not locally bounded around $\theta$. It follows that there exist $(\theta_{k})_{k\in\mathbb{N}}\subset\mathbb{R}^{q}$ and $(x_{k},\lambda_{k})\in\mathrm{Sol_{pd}}(\theta_{k})$, for all $k\in\mathbb{N}$, such that $\theta_{k}\xrightarrow[k\to\infty]{}\theta$ and $\lVert{(x_{k},\lambda_{k})}\rVert\xrightarrow[k\to\infty$$\infty$. Since $\mathrm{Sol_{p}}$ is locally bounded, we have that $(x_{k})_{k\in\mathbb{N}}$ is bounded and $\lVert{\lambda_{k}}\rVert\xrightarrow[k\to\infty$$\infty$. Passing to subsequences if necessary, we can assume that $\lVert{\lambda_{k}}\rVert\neq 0$ for all $k\in\mathbb{N}$, and $x_{k}\xrightarrow[k\to\infty]{}x$. By Lemma A.2(), we get $x\in\mathrm{Sol_{p}}(\theta)$. Let, for all $k\in\mathbb{N}$, $\alpha_{k}=\frac{\lambda_{k}}{\lVert{\lambda_{k}}\rVert}$, so that $\lVert{\alpha_{k}}\rVert=1$. Up to another subsequence, we can also assume $\alpha_{k}\xrightarrow[k\to\infty]{}\alpha$ such that $\lVert{\alpha}\rVert=1$. By definition of the set-valued map $\operatorname*{Mult}$, we know that, for every $k\in\mathbb{N}$,
$\exists\,v_{k}\in\operatorname*{conv\!}D_{F(\cdot,\theta_{k})}(x_{k}),\quad\operatorname{Jac}_{x}\!C(x_{k},\theta_{k})^{\!\mathsf{T}}\alpha_{k}=-\frac{v_{k}}{\lVert{\lambda_{k}}\rVert}.$ (4.3)
Using the fact that $(x,\theta)\mapsto D_{F(\cdot,\theta)}(x)$ is locally bounded (because $D_{F}$ is locally bounded), we have that $(v_{k})_{k\in\mathbb{N}}$ is bounded. Up to another subsequence if needed, we have that $v_{k}\xrightarrow[k\to\infty]{}v$. Taking the limit in (4.3), we can write
$\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\alpha=0.$ (4.4)
Furthermore, for all $k\in\mathbb{N}$, $\alpha_{k}\in K^{\circ}$ and $\left\langle\alpha_{k},C(x_{k},\theta_{k})\right\rangle=0$. Passing to the limit, we get
$\alpha\in K^{\circ}\;\text{ and }\;\left\langle\alpha,C(x,\theta)\right\rangle=0.$ (4.5)
Given Equations (4.4) and (4.5), the RCQ in Definition 4.1() imposes $\alpha=0$. This contradicts the fact that $\lVert{\alpha}\rVert=1$. Thus $\mathrm{Sol_{pd}}$ is locally bounded around $\theta$. ∎
@@
The last equality in (4.6) is due to the fact that, for all $(x,\lambda)\in\mathrm{Sol_{pd}}(\theta)$, one has by complementarity
$x\in\mathrm{Sol_{p}}(\theta),\;\lambda^{\!\mathsf{T}}C(x,\theta)=0.$
@@
To conclude, we will need the next Lemma which extends complementarity conditions for normal cones to time derivatives along smooth curves.
Lemma 4.8.
Let $K\subset\mathbb{R}^{m}$ be a closed convex cone and $z\colon\mathbb{R}\to K$, $\lambda\colon\mathbb{R}\to K^{\circ}$ be both continuously differentiable such that $\lambda(t)\in N_{K}(z(t))$ for all $t$. Then $\langle\dot{\lambda}(t),z(t)\rangle=\langle\lambda(t),\dot{z}(t)\rangle=0$ for all $t$.
We have $\dot{\lambda}(t)\in T_{K^{\circ}}(\lambda(t))$ for all $t$. We have for any $\lambda\in K^{\circ}$, using Lemma 4.2() and [42, Corollary 6.30]
$T_{K^{\circ}}(\lambda)=N_{K^{\circ}}(\lambda)^{\circ}=\{u\in K^{\circ\circ},\,\left\langle u,\lambda\right\rangle=0\}^{\circ}.$
Since $K^{\circ\circ}=K$, $z\in K$ and $\langle z,\lambda\rangle=0$, we deduce that for all $t$, $\langle z(t),\dot{\lambda}(t)\rangle\leq 0$. We have similarly for all $t$, $\langle\dot{z}(t),\lambda(t)\rangle\leq 0$ and, since for all $t$, $\langle z(t),\lambda(t)\rangle=0$, we obtain
$0=\frac{d}{dt}\langle z(t),\lambda(t)\rangle=\langle z(t),\dot{\lambda}(t)\rangle+\langle\dot{z}(t),\lambda(t)\rangle,$
from which the result follows. ∎
@@ 4.2 Parametric subdifferentiation of the value function
As discussed in the introduction, [42, Theorem 10.13] provides a formula for the subdifferential of $f$, which has a strong connection with the adjoint formula in (1.3). Let us make this explicit in the context of Problem (4.1). We will consider, only in this section, the extended formulation in order to motivate our construction. Let
@@
$(x,\theta)$ $\quad\mapsto\quad F(x,\theta)+\delta_{K}(C(x,\theta)),$
where, for $S\in\mathbb{R}^{\ell}$ with $\ell\in\mathbb{N}$, $\delta_{S}$ is the indicator function of $S$, i.e., $\delta_{S}(y)=0$ if $y\in S$ and $\delta_{S}(y)=\infty$ otherwise. Problem (4.1) is equivalent to unconstrained partial minimization of $\bar{F}$ so that this fits the context of [42, Theorem 10.13].
Under Assumption 4.4(), the RCQ holds jointly in $(x,\theta)$ as it is the case for the MFCQ in Remark 2.2(). Indeed, the first rows of the matrix $\operatorname{Jac}C(x,\theta)^{\!\mathsf{T}}$ are exactly those of $\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}$. Therefore, the chain rule [42, Theorem 10.6] applies to the indicator in the definition of $\bar{F}$. Note that the qualification condition is given in abstract form, it corresponds to Definition 4.1() using the fact that for any $z\in K$, $N_{K}(z)=\{\lambda\in K^{\circ},\,\left\langle\lambda,z\right\rangle=0\}$ [43, Corollary 23.5.4] and [42, Corollary 8.11, Definition 3.3] since $K$ is a closed convex cone.
Combining this with the sum rule [42, Exercise 10.10], we have, for all $(x,\theta)\in\operatorname*{gph}R$,
$\partial^{c}\bar{F}(x,\theta)$ $\subset\partial^{c}F(x,\theta)+\operatorname{Jac}C(x,\theta)^{\!\mathsf{T}}N_{K}(C(x,\theta)),$
where the normal cone is the subdifferential of the indicator in the definition of $\bar{F}$. In our case, the elements of the normal cones are given by the Lagrange multipliers $\lambda$ in $\mathrm{Sol_{pd}}(\theta)$; see Lemma 4.2(). Putting everything together, the parametric optimality condition, [42, Theorem 10.13], ensures that, under Assumptions 4.4(), for all $\theta\in\mathbb{R}^{q}$, $\partial^{c}f(\theta)$ is a subset of the following,
$D_{f}(\theta)\coloneq\operatorname*{conv\!}\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\partial^{c}F(x,\theta)+\operatorname{Jac}C(x,\theta)^{\!\mathsf{T}}\lambda\Big\}.$ (4.7)
Importantly, for all $\theta\in\mathbb{R}^{q}$ and for all $x\in\mathrm{Sol_{p}}(\theta)$, thanks to the parametric Fermat’s rule [42, Example 10.12], there exists $u\in\mathbb{R}^{q}$ such that $(0,u)\in\partial^{c}\bar{F}(x,\theta)$; hence $D_{f}(\theta)$ is nonempty.
However, as already stated earlier, $\partial^{c}f\neq D_{f}$ in general and the generalized adjoint formula produces artifacts; see Example 2.7(). We show in the coming section that, under Assumptions 4.4(), the set-valued map $D_{f}$ is a conservative field for $f$, providing a conservative adjoint formula.